OpenStudy (anonymous):

Find -3a^2 - b^3 + 3c^2 - 2b^3 if a = 2, b = -1, and c = 3. Numerical Answers Expected!

OpenStudy (anonymous):

Wow you have a lot of these!!

OpenStudy (anonymous):

Yea lol, I've been doing a lot of tests

OpenStudy (anonymous):

I needed to get my grade back up because it was an 89

OpenStudy (anonymous):

I'm trying to get access to my algebra final right now my teacher won't answer me Okay so I'm gonna solve this one on paper first because it takes a long time to type

OpenStudy (anonymous):

Ok haha

OpenStudy (anonymous):

I think I got it but what are your choices for this one

OpenStudy (anonymous):

It's not multiple choice it's a blank answer

OpenStudy (anonymous):

uggh alright well here's what I did -3(2)^2 -(-1)^3 + 3(3)^2 - 2(-1)^3 = -6^2 + 1 + 9^2 + 3^3 = 36 + 1 + 81 + 16 = 134 ??

OpenStudy (anonymous):

I got 12. :p

OpenStudy (anonymous):

welp that was wrong but owe well still got a 90 :P

OpenStudy (anonymous):

That's better than I would have gotten Yesterday I got an 85 on my practice algebra final Good job!!! :D:D